### v2

# Smooth polynomials with several prescribed coefficients

Abstract.
Let $\mathbb{F}_{q}[t]$ be the polynomial ring over the finite field $\mathbb{F}_{q}$ of $q$ elements. A polynomial in $\mathbb{F}_{q}[t]$ is called $m$-smooth (or $m$-friable) if all its irreducible factors are of degree at most $m$. In this paper, we investigate the distribution of $m$-smooth (or $m$-friable) polynomials with prescribed coefficients. Our technique is based on character sum estimates on smooth (friable) polynomials, Bourgains’s argument (2015) applied for polynomials by Ha (2016) and on double character sums on smooth (friable) polynomials.
Key words and phrases:
polynomial, finite field, exponential sums
2010 Mathematics Subject Classification:
11T06,11T55,11T23,11L07
1. Introduction
In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of the integers and some conditions on their digits in a given basis, see for example [2, 5, 6, 20, 21, 25].
In particular, Bourgain [2] investigated the distribution of primes with prescribed bits. Later, Swaenepoel [25] extended Bourgain’s result for arbitrary base. Namely, for a given positive integer $k$, consider its $g$-ary expansion
$k=\sum_{j\geq 0}k_{i}g^{i},\quad k_{i}\in\{0,1,\dots,g-1\}.$
For a given index-set ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ with #$\#{\mathcal{I}}\leq\delta(g)\cdot n$ and $0\in{\mathcal{I}}$, it follows from [2, 25], that the number of prime numbers with prescribed digits $\alpha_{i}\in\{0,1,\dots,g-1\}$ ($i\in{\mathcal{I}}$) with $\alpha_{0}\neq 0$ satisfies
#$\#\left\{p<g^{n},p_{i}=\alpha_{i},i\in{\mathcal{I}},\text{$p$ prime}\right\}\sim\frac{\pi(g^{n})}{g^{\#{\mathcal{I}}}},$
where $\pi(x)$ is the prime counting function.
Ha [12] investigated the same problem over function fields. Namely, let $\mathbb{F}_{q}$ be the finite field of $q$ elements, and consider the polynomial ring $\mathbb{F}_{q}[t]$ over $\mathbb{F}_{q}$. Let ${\mathcal{P}}(n)$ be the set of monic irreducible polynomials of degree $n$. It is well-known, that the number $\pi_{q}(n)=\#{\mathcal{P}}(n)$ of such irreducibly polynomials satisfies
(1) $\frac{q^{n}}{n}-2\frac{q^{n/2}}{n}\leq\pi_{q}(n)\leq\frac{q^{n}}{n},$
see e.g. [24, Lemma 4]. Improving earlier result on coefficients of irreducible polynomials [13, 24, 27], Ha showed that the number of irreducible polynomials with preassigned coefficients is proportional to $\pi_{q}(n)$. In particular, let ${\mathcal{M}}(n)$ be the set of monic polynomials of degree $n$ and for an index-set ${\mathcal{I}}\subset\{0,1\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) let ${\mathcal{J}}$ be the set of monic polynomials of degree $n$ with $i$th coefficient $\alpha_{i}$
(2) ${\mathcal{J}}=\left\{t^{n}+\sum_{i=0}^{n-1}c_{i}t^{i}\in{\mathcal{M}}(n):c_{i}=\alpha_{i},i\in{\mathcal{I}}\right\}.$
Write
$\delta=\frac{\#{\mathcal{I}}}{n},$
then for $\delta\leq 1/4$ we have
#$\#({\mathcal{P}}(n)\cap{\mathcal{J}})=\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}\left(1+O\left(\frac{1-\log_{q}\delta}{q^{1/\delta-4/(\delta+1)}}\right)\right)+O(q^{3n/4}),$
where
$\mathfrak{G}=\left\{\begin{array}[]{cl}1&\text{if $0\not\in{\mathcal{I}}$},\\ 1+\frac{1}{q-1}&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$},\\ 0&\text{if $0\in{\mathcal{I}}$ and $\alpha_{0}=0$}.\end{array}\right.$
In particular, we have the expected frequency of irreducible polynomials with prescribed coefficients, #$\#({\mathcal{P}}(n)\cap{\mathcal{J}})\sim\frac{\mathfrak{G}q^{n-\#{\mathcal{I}}}}{n}$, if $q\rightarrow\infty$ or $\delta=\#{\mathcal{I}}/n\rightarrow 0$.
In this paper, we investigate the distribution of smooth (or friable) polynomials with preassigned coefficients. For a given integer $m$, a polynomial is called $m$-smooth (or $m$-friable) if all its irreducible divisors are of degree at most $m$. Write
${\mathcal{S}}(n,m)=\{f\in{\mathcal{M}}(n):f\text{ is $m$-smooth}\}\quad\text{and}\quad\Psi(n,m)=\#{\mathcal{S}}(n,m).$
If we let
$u=\frac{n}{m},$
it is known, that the proportion of $m$-smooth polynomials of degree $n$ tends to Dickman’s $\rho$ function defined via the delayed differential equation
$\tau\rho^{\prime}(\tau)+\rho(\tau-1)=0$
for $\tau\geq 1$ with initial condition $\rho(\tau)=1$ for $\tau\in[0,1]$. It is decreasing, and de Bruijn proved [3] that it satisfies
$\rho(u)=\exp\left(-u\log\left(u\log u\right)+u+O\left(\frac{u\log\log u}{\log u}\right)\right)$
uniformly for $u\geq 3$.
Then Manstavic̆ius [19] proved that
(3) $\Psi(n,m)=q^{n}\rho\left(u\right)\left(1+O\left(\frac{u\log(u+1)}{m}\right)\right)$
holds in range $n\geq m\geq\sqrt{n\log n}$ (see also [4, 8, 23]). Recently, Gorodetsky [11] gave better approximation of $\Psi(n,m)$ with main term $q^{n}\rho(u-n/(2m^{2}))$ and improved error term.
The main result of this paper investigates the proportion of $m$-smooth polynomials with prescribed coefficients.
Theorem 1.
Let $40<\varepsilon<1/$ 4. Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and $\alpha_{i}\in\mathbb{F}_{q}$ ($i\in{\mathcal{I}}$) such that $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$. Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/10,\quad\text{that is,}\quad\#{\mathcal{I}}<n/10.$
Then we have
(4) #$\begin{split}\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})&=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)\\ &\qquad+O\left(\frac{q^{n}}{q^{\#I}}q^{-\frac{(1-\varepsilon)n+2m}{8}+2}+m^{2}n^{1/2}q^{(7n+m+4)/8}\right),\end{split}$
holds for some positive $C>0$ in range $n\geq m\geq\sqrt{n\log n}$, where the implied constant may depend on $\varepsilon$.
In particular, the number of $m$-smooth polynomials with prescribed coefficients has the expected frequency if $q\rightarrow\infty$ or $\delta\rightarrow 0$.
Corollary 2.
Assume, that
  • for fixed $n$, $u$ and $\delta$, we have $q\rightarrow\infty$; or
  • for fixed $q$, we have $n\rightarrow\infty$ and $\delta=o(n)$.
Then
(5) #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})\sim q^{-\#I}\Psi(n,m).$
If $\alpha_{0}=0$, then #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ deviates from the expected frequency. For example, if $I=\{0,1,\dots,r-1\}$ with $\alpha_{i}=0$ for $i\in{\mathcal{I}}$, then
${\mathcal{S}}(n,m)\cap{\mathcal{J}}=\{t^{n}f:f\in{\mathcal{S}}(n-r,m)\}$
and thus by (3),
(6) #$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})=\Psi(n-r,m)\sim\rho\left(\frac{n-r}{m}\right)q^{n-r}$
in contrast to (4).
In order to handle the general case write
$I=\{i_{0}<\dots<i_{\nu}<\dots\}$
with $\alpha_{i_{0}}=\dots=\alpha_{i_{\nu-1}}=0\neq\alpha_{i_{\nu}}$ and define
$\Lambda_{0}(n,m,I,\nu)=\frac{-1}{m}\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<m/8\end{subarray}}q^{j-i_{j}}\rho^{\prime}\left(\frac{n-i_{j}}{m}\right)$
and
$\Lambda_{1}(n,m,I,\nu)=\left\{\begin{array}[]{cl}-\frac{1}{m}\frac{q^{\nu-i_{\nu}}}{q-1}\rho^{\prime}\left(\frac{n-i_{\nu}}{m}\right)&\text{if $i_{\nu}<m/8$},\\ 0&\text{otherwise}.\end{array}\right.$
Then we have
Theorem 3.
Let ${\mathcal{I}}\subset\{0,1,\dots,n-1\}$ and choose $\alpha_{i}\in\mathbb{F}_{q}$ for each $i\in{\mathcal{I}}$. Write $\delta=\#{\mathcal{I}}/n$ and assume that
$\delta<1/10,\quad\text{that is,}\quad\#{\mathcal{I}}<n/10,$
then we have
#$\#({\mathcal{S}}(n,m)\cap{\mathcal{J}})$ $=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q^{i_{0}+1}}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)$
$\qquad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)+O\left(q^{-\frac{(1-\varepsilon)n+2m}{8}+2}e^{\varepsilon(n-m)}\right)\right)$
$\qquad+O\left(m^{2}n^{1/2}q^{(7n+m+4)/8}\right),$
holds for some positive $C>0$ in range $n\geq m\geq\sqrt{n\log n}$.
By Theorem 3, the distribution of $m$-smooth polynomials only depends on the position of the prescribed zero coefficients.
We remark, if $0\in{\mathcal{I}}$, then we can apply either Theorem 1 if $\alpha_{0}\neq 0$ or (6) if $\alpha_{0}=0$. In the case $0\not\in{\mathcal{I}}$, Theorem 3 yields the expected frequency if $q\rightarrow\infty$ or $\delta$ is small enough in terms of $u$.
Corollary 4.
Assume, that $0\not\in{\mathcal{I}}$ and
  • if $q\rightarrow\infty$ for fixed $n$ and $u$; or
  • if $n\rightarrow\infty$ with $\delta=o\left((u\log u)^{9/8}\right)$ and fixed $q$.
Then (5) holds.
Theorems 1 and 3 give asymptotics to the number smooth polynomials with several prescribed coefficients. Recently, Kuttner and Wang [18], using different techniques, have considered the exact number of smooth polynomials of degree $n$ with given trace, that is, when =${\mathcal{I}}=\{n-1\}$.
In a different perspective, Gorodetsky [9] also investigated the number of smooth polynomials with three prescribed coefficients in characteristic 2 and 5 and reveal, after normalization, a cyclic property of their number in terms of $n$.
Hauck and Shparlinski [14] has also considered similar problems over $\mathbb{Z}$ by proving the existence of smooth integers (that is, integers without large prime factor) with few nonzero binary digits.
The proof of Theorem 3 uses the circles method. The major arc analysis is based on character sum estimates on smooth polynomials, see Section 4, and Bourgain’s argument [2], applied for polynomials over finite fields by Ha [12], see Section 6.
The minor arc analysis is based on double exponential sums and the good factorisation properties of smooth polynomials, see Section 5.2.
2. Outline of the proof
For $f\in\mathbb{F}_{q}[t]$, write
(7) $|f|=q^{\deg f}$
with the convention $|0|=0$. One can extend the absolute value ||$\cdot$|| ⋅ | to $\mathbb{F}_{q}(t)$ in the natural way by $|f/g|=q^{\deg f-\deg g}$.
Let $\mathbb{K}_{\infty}$ be the set of formal power series
$\mathbb{K}_{\infty}=\left\{\xi=\sum_{i\leq k}x_{i}t^{i},k\in\mathbb{Z},x_{i}\in\mathbb{F}_{q}\right\},$
which is the completion of $\mathbb{F}_{q}(t)$ at the prime associated to the $(1/t)$-adic valuation.
We extend the norm (7) to $\mathbb{K}_{\infty}$ by
$|\xi|=q^{k},$
where $k$ is the larges index so that $x_{k}\neq 0$.
Also, for $\xi\in\mathbb{K}_{\infty}$ we define
$\|\xi\|=\min_{f\in\mathbb{F}_{q}[t]}|\xi-f|.$
Write
$\mathbf{e}(\xi)=\exp\left(\frac{2\pi i}{p}\mathrm{tr}_{\mathbb{F}_{q}}(x_{-1})\right),$
where $\mathrm{tr}_{\mathbb{F}_{q}}$ is the (absolute) trace of $\mathbb{F}_{q}$ and $p$ is the characteristic of $\mathbb{F}_{q}$.
We define $\mathbf{T}$ by $\{\xi\in\mathbb{K}_{\infty}:|\xi|<1\}$, and fix an additive Haar measure, normalized so that $\int_{\mathbf{T}}\mathrm{d}\xi=1$. Then we have for a polynomial $f\in\mathbb{F}_{q}[t]$,
(8) $\int_{\mathbf{T}}\mathbf{e}(f\xi)\mathrm{d}\xi=\left\{\begin{array}[]{cl}1&\text{if }f=0,\\ 0&\text{if }f\neq 0,\end{array}\right.$
see [16, Theorem 3.5].
For given integers $m\leq n$, write
$S(\xi;n,m)=\sum_{f\in{\mathcal{S}}(n,m)}\mathbf{e}(f\xi).$
For a nonempty subset ${\mathcal{I}}\subset\{0,1,\dots n-1\}$ fix $\alpha_{i}\in\mathbb{F}_{q}[t]$ for $i\in{\mathcal{I}}$. Let ${\mathcal{J}}$ defined by (2) and write
$S_{{\mathcal{J}}}(\xi)=\sum_{f\in{\mathcal{J}}}\mathbf{e}(f\xi).$
By (8), we have
(9) $\int_{\mathbf{T}}S(\xi;n,m)\overline{S_{{\mathcal{J}}}(\xi)}\mathrm{d}\xi=\sum_{f\in{\mathcal{S}}(n,m)}\sum_{g\in{\mathcal{J}}}\int_{\mathbf{T}}\mathbf{e}((f-g)\xi)\mathrm{d}\xi=\#({\mathcal{S}}(n,m)\cap{\mathcal{J}}).$
We investigate this integral by the circle method. To do so, let $\kappa,\ell$ be two positive parameters to be fixed later such that
$\ell+\kappa<n$
and define the major and minor arcs by
(10) ${\mathfrak{M}}=\bigcup_{|a|<|g|\leq q^{\kappa}}\left\{\xi\in{\mathbf{T}}:\left|\xi-\frac{a}{g}\right|<q^{-n+\ell}\right\}\quad\text{and}\quad{\mathfrak{m}}={\mathbf{T}}\setminus{\mathfrak{M}}.$
We give approximation of $S(\xi;n,m)$ on major arcs in Section 5.1 using character sum estimates on smooth polynomials introduced in Section 4. The minor arc estimates of $S(\xi;n,m)$ are based on double exponential sums and the good factorisation properties of smooth polynomials, see Section 5.2.
The analysis of $S_{{\mathcal{J}}}(\xi)$ is based on Bourgain’s argument [2], applied for polynomials over finite fields by Ha [12], see Section 6.
3. Preliminaries
3.1. Notations
For given functions $F$ and $G$, the notations $F\ll G$, $G\gg F$ and $F=O(G)$ are all equivalent to the statement that the inequality $|F|\leq c|G|$ holds with some constant $c>0$. Throughout the paper, any implied constants in symbols O, $\ll$ and $\gg$ are absolute unless specified otherwise. For a set ${\mathcal{A}}$ we denote the characteristic function of ${\mathcal{A}}$ by $\mathbf{1}({\mathcal{A}})$.
We let ${\mathcal{M}}=\bigcup_{n}{\mathcal{M}}(n)$ and ${\mathcal{P}}=\bigcup_{n}{\mathcal{P}}(n)$ to be the sets of all monic and all monic irreducible polynomials respectively.
3.2. Arithmetically distributed relations
For a monic polynomial $f=t^{n}+c_{n-1}t^{n-1}+\dots+c_{0}$, we call $c_{n-1},\dots,c_{n-\ell}$ the first $\ell$ coefficients of $f$. Define $f^{*}=t^{\deg f}f(1/t)$. For $\ell\geq 0$ and $g\in\mathbb{F}_{q}[t]$, define the arithmetically distributed relation ${\mathcal{R}}_{\ell,g}$ as
$f\equiv h\mod{\mathcal{R}}_{\ell,g}\quad\text{if}\quad f\equiv h\mod g\quad\text{and}\quad f^{*}\equiv h^{*}\mod t^{\ell+1},$
that is, if $f$ and $h$ are congruent modulo $g$ and their first $\ell$ coefficients are the same. Clearly, an arithmetically distributed relation ${\mathcal{R}}_{\ell,g}$ is an equivalence relation.
We also write ${\mathcal{R}}_{\ell}={\mathcal{R}}_{\ell,1}$ and ${\mathcal{R}}_{g}={\mathcal{R}}_{0,g}$.
Observe, that ${\mathcal{R}}_{\ell}$ on fixed degree polynomials captures the polynomials on intervals, that is, for polynomials $f,h\in{\mathcal{M}}(n)$,
$f\equiv h\mod{\mathcal{R}}_{\ell}\quad\text{if and only if}\quad|f-h|<q^{n-\ell}.$
The polynomial $f$ is invertible modulo ${\mathcal{R}}_{\ell,g}$ if $\gcd(f,g)=1$. These invertible polynomials form a group, denoted by $G_{\ell,g}=\left({\mathcal{M}}/{\mathcal{R}}_{\ell,g}\right)^{\times}$. Write $G_{\ell}=\left({\mathcal{M}}/{\mathcal{R}}_{\ell}\right)^{\times}$ and $G_{g}=\left({\mathcal{M}}/{\mathcal{R}}_{g}\right)^{\times}$. Then we have
(11) $G_{\ell,g}\cong G_{\ell}\times G_{g},$
see [15, Theorem 8.6] and [17, Lemma 1.1]. We also have #$\#G_{\ell}=q^{\ell}$ and
#$\#G_{g}=\#(\mathbb{F}_{q}[t]/(g))^{\times}=\Phi(g).$
We have the following lower bound on $\Phi(g)$, see [12, Lemma 2.3],
(12) $\Phi(H)\gg\frac{|g|}{\log_{q}(\deg g)+1}\quad\text{for }g\neq 0.$
Let $\chi$ be a character modulo an arithmetically distributed relation ${\mathcal{R}}_{\ell,g}$, and let $L(s,\chi)$ be the associated $L$-function,
$L(s,\chi)=\sum_{f\in{\mathcal{M}}}\frac{\chi(f)}{|f|^{s}}=\prod_{\begin{subarray}{c}\omega\nmid g\\ \omega\in{\mathcal{P}}\end{subarray}}\left(1-\frac{\chi(\omega)}{|\omega|^{s}}\right)^{-1},\quad\mathrm{Re}(s)>1.$
The following result, see [17, Theorem 1.3], follows from the Riemann hypothesis for the $L$-functions of algebraic curves over finite fields proved by Weil [28].
Lemma 5.
If $\chi$ is a non-trivial multiplicative character modulo ${\mathcal{R}}_{\ell,g}$, then there exists a polynomial $P_{\chi}$ of degree at most $\ell-1+\deg g$ and of form
$P_{\chi}(z)=\prod_{i}\left(1-\alpha_{i}z\right)$
with $L(s,\chi)=P_{\chi}(q^{-s})$ and $|\alpha_{i}|=\sqrt{q}$ with at most one exceptional $\alpha_{i}$ with $|\alpha_{i}|=1$.
Using Lemma 5, on can derive the following character sum estimates along irreducible polynomials, see [17] for more details.
Lemma 6.
Let $\chi$ be a non-trivial character modulo ${\mathcal{R}}_{\ell,g}$. Then
$\left|\sum_{\omega\in{\mathcal{P}}(n)}\chi(\omega)\right|\leq\frac{1}{n}(\ell-1+\deg g)q^{n/2}.$
3.3. Exponential sums
We have, see [16, Theorem 3.7].
Lemma 7.
Suppose $|\xi|<1$. Then
$\sum_{f\in{\mathcal{M}}(n)}\mathbf{e}(f\xi)=\left\{\begin{array}[]{lc}q^{n}\mathbf{e}(t^{n}\xi)&\text{if }|\xi|<q^{n},\\ 0&\text{otherwise}.\end{array}\right.$
We have, see [16, Equation (4.6)].
Lemma 8.
If $a,b$ are invertible modulo ${\mathcal{R}}_{\ell,g}$, then
$\frac{1}{q^{\ell}\Phi(g)}\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\bar{\chi}(a)\chi(b)=\left\{\begin{array}[]{ll}1&\text{if }a\equiv b\mod{\mathcal{R}}_{\ell,g},\\ 0&\text{otherwise}.\end{array}\right.$
For the following result see [16, Theorem 6.1].
Lemma 9.
For polynomials $f,g$, we have
$\sideset{}{{}^{*}}{\sum}_{a\bmod g}\mathbf{e}(af/g)=\sum_{d\mid\gcd(f,g)}|d|\mu(g/d).$
We can prove the following results on polynomial Gauss sums.
Lemma 10.
For polynomials $b,g$, we have
$\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a\bmod{\mathcal{R}}_{\ell,g}\\ a\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\chi(a)\mathbf{e}(a/g)\right|^{2}=q^{\ell}\Phi(g)^{2}$
Proof.
We have by Lemma 8, that
$\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a\bmod{\mathcal{R}}_{\ell,g}\\ a\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\chi(a)\mathbf{e}(a/g)\right|^{2}$
$=\sum_{\chi\bmod{\mathcal{R}}_{\ell,g}}\left(\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{1}\bmod{\mathcal{R}}_{\ell,g}\\ a_{1}\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\chi(a_{1})\mathbf{e}(a_{1}/g)\right)\left(\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{2}\bmod{\mathcal{R}}_{\ell,g}\\ a_{2}\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}(a_{2})\mathbf{e}(-a_{2}/g)\right)$
[…]
$=q^{\ell}\Phi(g)\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}a_{1}\equiv a_{2}\bmod{\mathcal{R}}_{g}\\ a_{1}\equiv a_{2}\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}((a-b)/g)=q^{\ell}\Phi(g)^{2}.$
For the following estimate see [22, Lemma 4].
Lemma 11.
Let
$\xi=\frac{a}{g}+\gamma\quad\text{with}\quad|\gamma|\leq\frac{1}{|g|^{2}}\quad\text{and}\quad(a,g)=1.$
Then we have
$\sum_{f\in{\mathcal{M}}(k)}\left|\sum_{h\in{\mathcal{M}}(j)}\mathbf{e}(\xi fh)\right|\ll\frac{q^{k+j}}{|g|}+q^{k}+|g|.$
3.4. Dickman’s $\rho$ function
We need the following result on Dickman’s $\rho$ function which follows directly from [26, Lemma III. 5.11] and [26, Corollary III. 5.15].
Lemma 12.
Uniformly over $0\leq v\leq u$, we have
$\rho(u-v)\ll(u\log u)^{v}\rho(u).$
For the following result see [7, Lemma 4.1].
Lemma 13.
For any integer $k\geq 0$ we have
$(-1)^{k}\rho^{(k)}(u)>0\quad\text{for $u>1$}$
and
$\rho^{(k)}(u)\sim(-1)^{k}\rho(u)(\log u)^{k}\quad\text{as }u\rightarrow\infty.$
3.5. Further preliminaries
The following lemma of Hayes [16, Theorem 4.3] is an analogue of a well-known result of Dirichlet in the theory of Diophantine approximation.
Lemma 14.
For any integer $n$ and $\xi\in\mathbf{T}$, there is a pair of coprime polynomials $a,g\in\mathbb{F}_{q}[t]$ with $g$ monic, $|a|<|g|\leq q^{n/2}$ such that
$\left|\xi-\frac{a}{g}\right|<\frac{1}{|g|q^{n/2}}.$
Let $\tau(f)$ denote the number of divisors of $f$. We have the following bound on $\tau(f)$.
Lemma 15.
For any $\varepsilon>0$ there is a constant $c(\varepsilon)>0$ such that
$\tau(f)\leq c(\varepsilon)|f|^{(2+\varepsilon)/\log\deg f}.$
The proof of this lemma is analogue to the integer case, however, we include its proof for the sake for completeness.
Proof.
Write
$f=g_{1}^{e_{1}}\dots g_{k}^{e_{k}}$
with pairwise distinct irreducible polynomials $g_{1},\dots,g_{k}$. Then
$\tau(f)=\prod_{i=1}^{k}(e_{i}+1).$
For any $\lambda>0$ we have
$\frac{\tau(f)}{\exp(\lambda\deg f)}=\prod_{i=1}^{k}\frac{e_{i}+1}{\exp(\lambda e_{i}\deg g_{i})}\leq\prod_{i=1}^{k}\frac{e_{i}+1}{1+\lambda e_{i}\deg g_{i}}.$
For $\deg g_{i}>1/\lambda$ we have
$\frac{e_{i}+1}{1+\lambda e_{i}\deg g_{i}}<1.$
Thus
$\frac{\tau(f)}{\exp(\lambda\deg f)}\leq\prod_{\deg g_{i}\leq 1/\lambda}\frac{e_{i}+1}{1+\lambda e_{i}\deg g_{i}}.$
As there are at most $q^{1/\lambda}$ irreducible polynomial of degree at most $1/\lambda$, we get
(13) $\frac{\tau(f)}{\exp(\lambda\deg f)}\leq\left(\frac{1}{\lambda}\right)^{q^{1/\lambda}}=\exp\left(q^{1/\lambda}\log(1/\lambda)\right).$
Choosing
$\lambda=\frac{2\log q}{\log\deg f},$
we get that the logarithm of (13) is
$\log\frac{\tau(f)}{\exp\left(\frac{2\log q\deg f}{\log\deg f}\right)}\leq(\deg f)^{1/2}\log\frac{\log\deg f}{2\log q}\leq\varepsilon\frac{\log q\deg f}{\log\deg f}$
if $\deg f$ is large enough in terms of $\varepsilon$. ∎
4. Distribution of smooth polynomials
The purpose of this section is to prove character sum estimates along smooth polynomial. The proof is based on the method of Bhowmick, Lê and Liu [1], see also [10]. We consider both cases of $q$ or $n$ are fixed in contrast to [1, 10].
Lemma 16.
Let $\chi\neq\chi_{0}$ be a character modulo ${\mathcal{R}}_{\ell,g}$ and $n>\ell-1+\deg g$.
  1. (i)
    We have
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\frac{n}{2}+\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}}e^{13q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}$
    if $\ell+\deg g$ is large enough in terms of $q$. One can choose
    $\ell+\deg g>10^{4}\quad\text{and}\quad\frac{\log(\ell+1+\deg g)}{\log\log(\ell+1+\deg g)}>\log q.$
  2. (ii)
    Let $40<\varepsilon<1/$ 4, then we have
    $\left|\sum_{f\in S(n,m)}\chi(f)\right|\leq q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+1+\deg g)}$
    if $q$ is large enough in terms of $\varepsilon$. One can choose $q>(1+1/\varepsilon)^{1/\varepsilon}$.
Proof.
Write
$T(z,\chi;m)=\prod_{\omega\in{\mathcal{P}},\deg\omega\leq m}\left(1-\chi(\omega)z^{\deg\omega}\right)^{-1}.$
For $z\in\mathbb{C}$ with $|z|<1$, we have
$T(z,\chi;m)=\sum_{r=0}^{\infty}S(\chi;r,m)z^{r},\quad S(\chi;r,m)=\sum_{f\in{\mathcal{S}}(r,m)}\chi(f).$
The logarithm of $T(z,\chi;m)$ is given by
(14) $\log T(z,\chi;m)=\sum_{\omega\in{\mathcal{P}},\deg\omega\leq m}\chi(\omega)z^{\deg\omega}+\sum_{k\geq 2}k^{-1}\sum_{\omega\in{\mathcal{P}},\deg\omega\leq m}\chi(\omega)^{k}z^{k\deg\omega}.$
By Lemma 6, we have
(15) $\left|\sum_{\omega\in{\mathcal{P}}(r)}\chi(\omega)\right|\leq\frac{\ell-1+\deg g}{r}q^{r/2}.$
We also have the trivial bound
(16) $\left|\sum_{\omega\in{\mathcal{P}}(r)}\chi(\omega)\right|\leq\frac{q^{r}}{r}$
by (1).
Let $R$ be a positive integer to be fixed later.
For for $r\geq R$ we use (15) in (14), while $r<R$ we use the trivial bound (16). Namely, (15) yields that for $r\geq R$, we have that the coefficient of $z^{r}$ in (14) is bounded in absolute value by
(17) $\frac{\ell-1+\deg g}{r}q^{r/2}+\sum_{\begin{subarray}{c}k\mid r\\ k\geq 2\end{subarray}}k^{-1}\#{\mathcal{P}}(r/k)\leq\frac{\ell+1+\deg g}{r}q^{r/2}.$
For $r<R$, the trivial bound (16) yields, that the coefficient of $z^{r}$ in (14) is bounded in absolute value by
(18) $\frac{q^{r}}{r}+\sum_{\begin{subarray}{c}k\mid r\\ k\geq 2\end{subarray}}k^{-1}\#{\mathcal{P}}(r/k)\leq\frac{2}{r}q^{r}.$
Combining (17) and (18), we get
(19) $\log T(z,\chi;m)\leq\sum_{1\leq r<R}\frac{2}{r}(q|z|)^{r}+\sum_{r\geq R}\frac{\ell+1+\deg g}{r}(q^{1/2}|z|)^{r}.$
In order to prove (i), put
$R=\left\lfloor 2\frac{\log(\ell+1+\deg g)}{\log q}\right\rfloor+1.$
With this choice, we have
(20) $q^{(R-1)/2}\leq\ell+1+\deg g\leq q^{R/2}.$
Also, let
(21) $0<\delta<\min\left\{\frac{1}{\log q},\frac{1}{4}\right\}$
to be chosen later. With this choice, put $Z=q^{-1/2-\delta}$ and let $|z|=Z$.
Then for the first term of (19) we have
(22) $\sum_{1\leq r<R}\frac{2}{r}(q|z|)^{r}\leq 2(qZ)^{R-1}\sum_{r\geq 0}(qZ)^{-r}\leq\frac{2(\ell+1+\deg g)^{2}Z^{R-1}}{1-(qZ)^{-1}}\leq 13(\ell+1+\deg g)^{2}Z^{R-1}$
using $|qZ|\geq q^{1/4}\geq 2^{1/4}$.
Also,
(23) $\sum_{r\geq R}\frac{\ell+1+\deg g}{r}(q^{1/2}Z)^{r}\leq\frac{\ell+1+\deg g}{R}\cdot\frac{(q^{1/2}Z)^{R}}{1-q^{1/2}Z}.$
Then combining (19), (22) and (23), we have by (20) that
$\log T(z,\chi;m)$ $\leq 13(\ell+1+\deg g)^{2}q^{-(1/2+\delta)(R-1)}+\frac{\ell+1+\deg g}{R}\cdot\frac{q^{-\delta R}}{1-q^{-\delta}}$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{R(1-q^{-\delta})}\right).$
As $\delta<1/\log q$, we have $1-q^{-\delta}\geq\frac{1}{2}\delta\log q$, thus
$\log T(z,\chi;m)$ $\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{R\delta\log q}\right)$
$\leq(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right).$
Let $C_{Z}$ denote the circle centered at 0 with radius $Z$. Then, by the Cauchy’s integral formula, we have
$|S(\chi,n,m)|$ $=\left|\frac{1}{2\pi i}\int_{C_{Z}}T(z,\chi;m)z^{-n-1}\mathrm{d}z\right|$
$\leq\max_{C_{Z}}|T(z,\chi;m)|Z^{-n}$
$\leq\exp\left((\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)\right)q^{(1/2+\delta)n}.$
Put
$\delta=\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}$
and observe that it satisfies (21) by assumption. If $\ell+1+\deg g>10^{4}$, then $\delta\leq 1/4$, and also, $\delta\leq\frac{1}{\log q}$. Then
$(\ell+1+\deg g)^{(1-2\delta)}\left(13q^{1/2}+\frac{1}{\delta\log(\ell+1+\deg g)}\right)$
$=$ $\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}\left(13q^{1/2}+\frac{1}{\log\log(\ell+1+\deg g)}\right)$
$\leq$ $13q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}$
which proves (i).
In order to prove (ii), we choose $R=1$, that is, in (19) the first term vanishes and
$\log T(z,\chi;m)\leq\sum_{r\geq 1}\frac{\ell+1+\deg g}{r}(q^{1/2}|z|)^{r}.$
Choose $Z=q^{-1/2-\varepsilon}$. For $|z|=Z$, we have
$\log T(z,\chi;m)\leq\sum_{r\geq 1}\frac{\ell+1+\deg g}{r}(q^{-\varepsilon})^{r}\leq(\ell+1+\deg g)\cdot\frac{q^{-\varepsilon}}{1-q^{-\varepsilon}}\leq\varepsilon(\ell+1+\deg g)$
if $q$ is large enough in terms of $\varepsilon$.
Then, by the Cauchy’s integral formula, we have
$|S(\chi,n,m)|$ $=\left|\frac{1}{2\pi i}\int_{C_{Z}}T(z,\chi;m)z^{-n-1}\mathrm{d}z\right|$
$\leq\max_{C_{Z}}|T(z,\chi;m)|Z^{-n}$
$\leq\exp\left(\varepsilon(\ell+1+\deg g)\right)q^{(1/2+\varepsilon)n},$
which proves (ii). ∎
The following result unifies the two bounds of Lemma 16.
Corollary 17.
Let $\chi\neq\chi_{0}$ be a character modulo ${\mathcal{R}}_{\ell,g}$ and $n\geq\ell+\deg g$. For all $40<\varepsilon<1/$ 4, we have
$\sum_{f\in S(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)},$
where the implied constant may depend on $\varepsilon$.
Proof.
We can assume, that $q^{n}$ is large enough in terms of $\varepsilon$.
If $q>(1+1/\varepsilon)^{1/\varepsilon}$, then we apply (ii) of Lemma 16 directly. Otherwise,
$e^{13q\frac{\ell+1+\deg g}{\log^{2}(\ell+1+\deg g)}}\ll e^{\varepsilon(\ell+\deg g)}$
and if $n$ is large enough, then
$q^{\frac{\log\log(\ell+1+\deg g)}{\log(\ell+1+\deg g)}}\ll q^{\varepsilon n}.$
As an application, we prove the following result on the distribution of smooth polynomials in intervals.
Lemma 18.
Let $40<\varepsilon<1/$ 4. For $1\leq\ell\leq n$ we have for $b\in\mathbb{F}_{q}[t]$ that
#$\#\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}=\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right),$
where the implied constant may depend on $\varepsilon$.
Proof.
We have
(24) #$\#\{f\in{\mathcal{S}}(n,m):f\equiv b\mod{\mathcal{R}}_{\ell}\}$ $=\frac{1}{q^{\ell}}\sum_{\chi\bmod{\mathcal{R}}_{\ell}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f).$
The contribution of the principal character $\chi_{0}$ is
$\frac{1}{q^{\ell}}\overline{\chi_{0}(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi_{0}(f)=\frac{1}{q^{\ell}}\Psi(n,m).$
By Corollary 17, the contribution of the non-principal characters is
$\frac{1}{q^{\ell}}\sum_{\chi\neq\chi_{0}}\overline{\chi(b)}\sum_{f\in{\mathcal{S}}(n,m)}\chi(f)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}.$
5. Analysis of ${\mathcal{S}}(\xi;n,m)$
In this section we investigate ${\mathcal{S}}(\xi;n,m)$ on major and minor arcs.
5.1. Major arc estimates
Let $\Psi_{g}(n,m)$ denote the number of $m$-smooth polynomials of degree $n$ which are coprime to $g$,
$\Psi_{g}(n,m)=\#\{f\in{\mathcal{S}}(n,m):\gcd(f,g)=1\}.$
We split the summation over $f$ in ${\mathcal{S}}(\xi;n,m)$ into residue classes modulo ${\mathcal{R}}_{\ell}$
${\mathcal{S}}(\xi;n,m)=\sum_{b\bmod{\mathcal{R}}_{\ell}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(f\xi\right)$
and investigate each term on the major arcs.
Lemma 19.
Let $40<\varepsilon<1/$ 4. For $a,b\in\mathbb{F}_{q}[t]$, $\gcd(a,g)=1$, $\ell+\deg g<n$, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)$ $=\sum_{d\mid g}\frac{\mu(g/d)}{q^{\ell}\Phi(g/d)}\Psi_{g/d}(n-\deg d,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}|g|^{1/2}\right),$
where the implied constant may depend on $\varepsilon$.
Proof.
We have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)=\sum_{d\mid g}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\\ \gcd(f,g/d)=1\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g/d}\right).$
For a given $d$ we have by Lemma 8 that
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\\ \gcd(f,g/d)=1\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g/d}\right)$ $=\frac{1}{q^{\ell}\Phi(g/d)}\sum_{\chi\bmod{\mathcal{R}}_{g/d,\ell}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\end{subarray}}\chi(f)\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}(c)\mathbf{e}\left(\frac{ac}{g/d}\right),$
where in the innermost sum $c\in{\mathcal{M}}(n-\deg d)$ runs over all resuced residue calsses modulo ${\mathcal{R}}_{g/d,\ell}$ such that $c\equiv b\mod{\mathcal{R}}_{\ell}$. As $n-\deg d>\deg(g/d)+\ell$, by (11), the polynomial $c$ runs thorough all residue classes modulo $g/d$. Thus by Lemma 9, the contributions of the principal characters $\chi_{0}^{(g/d,\ell)}$ modulo ${\mathcal{R}}_{g/d,\ell}$ is
$\sum_{d\mid g}\frac{1}{q^{\ell}\Phi(g/d)}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\end{subarray}}\chi_{0}^{(g/d,\ell)}(f)\sideset{}{{}^{*}}{\sum}_{c\bmod g/d}\mathbf{e}\left(\frac{ac}{g/d}\right)=\sum_{d\mid g}\frac{\mu(g/d)}{q^{\ell}\Phi(g/d)}\Psi_{g/d}(n-\deg d,m).$
By (11) and Lemma 17, the contribution of the non-principal characters are
$\frac{1}{q^{\ell}\Phi(g/d)}\left|\sum_{\begin{subarray}{c}\chi^{(g/d,\ell)}\bmod{\mathcal{R}}_{g/d,\ell}\\ \chi^{(g/d,\ell)}\neq\chi^{(g/d,\ell)}_{0}\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\end{subarray}}\chi^{(g/d,\ell)}(f)\sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}^{(g/d,\ell)}(a)\mathbf{e}\left(\frac{ac}{g/d}\right)\right|$
$\leq\frac{1}{q^{\ell}\Phi(g/d)}\sum_{\begin{subarray}{c}\chi^{(g/d,\ell)}\bmod{\mathcal{R}}_{g/d,\ell}\\ \chi^{(g/d,\ell)}\neq\chi^{(g/d,\ell)}_{0}\end{subarray}}\left|\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n-\deg d,m)\end{subarray}}\chi^{(g/d,\ell)}(f)\right|\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}^{(g/d,\ell)}(a)\mathbf{e}\left(\frac{ac}{g/d}\right)\right|$
$\ll q^{\left(\frac{1}{2}+\varepsilon\right)(n-\deg d)}e^{\varepsilon(\ell+\deg g)}\frac{1}{q^{\ell}\Phi(g/d)}\sum_{\begin{subarray}{c}\chi^{(g/d,\ell)}\bmod{\mathcal{R}}_{g/d,\ell}\end{subarray}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}^{(g/d,\ell)}(a)\mathbf{e}\left(\frac{ac}{g/d}\right)\right|.$
By the Cauchy-Schwartz inequality and Lemma 10 we have
$\left(\ \sum_{\begin{subarray}{c}\chi^{(g/d,\ell)}\bmod{\mathcal{R}}_{g/d,\ell}\end{subarray}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}^{(g/d,\ell)}(a)\mathbf{e}\left(\frac{ac}{g/d}\right)\right|\right)^{2}$
$\leq q^{\ell}\Phi(g/d)\sum_{\begin{subarray}{c}\chi^{(g/d,\ell)}\bmod{\mathcal{R}}_{g/d,\ell}\end{subarray}}\left|\ \sideset{}{{}^{*}}{\sum}_{\begin{subarray}{c}c\bmod{\mathcal{R}}_{g/d,\ell}\\ c\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\bar{\chi}^{(g/d,\ell)}(a)\mathbf{e}\left(\frac{ac}{g/d}\right)\right|^{2}$
$\leq q^{2\ell}\Phi(g/d)^{3}.$
Thus, summing over $d$ we get, that the error term is
$\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}\sum_{d\mid g}q^{-\left(\frac{1}{2}+\varepsilon\right)\deg d}\Phi(g/d)^{1/2}\ll q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon(\ell+\deg g)}|g|^{1/2}\sum_{d\mid g}\frac{1}{|d|^{1+\varepsilon}}$
As
$\sum_{d\mid g}\frac{1}{|d|^{1+\varepsilon}}=\sum_{i\leq\deg g}\frac{1}{q^{i(1+\varepsilon)}}\sum_{\begin{subarray}{c}d\mid g\\ \deg d=i\end{subarray}}1\leq\sum_{i\leq\deg g}\frac{1}{q^{\varepsilon i}}\leq\sum_{i=1}^{\infty}\frac{1}{q^{\varepsilon i}}\ll 1,$
the result follows. ∎
The following result estimates the main term of Lemma 19 for polynomials $g$ of form $g=t^{k}$.
Lemma 20.
Let $40<\varepsilon<1/$ 4. For $b\in\mathbb{F}_{q}[t]$, $\gcd(b,t)=1$, $1\leq k<n-\ell$, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{bf}{t^{k}}\right)$ $=-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)$
$\quad+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{k}{2}}e^{\varepsilon k}\right),$
where the implied constant may depend on $\varepsilon$.
Proof.
By Lemma 19, the main term is
$\sum_{i=0}^{k}\frac{\mu(t^{k-i})}{q^{\ell}\Phi(t^{k-i})}\Psi_{t^{k-i}}(n-i,m)$
$=\frac{1}{q^{\ell}}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi_{t}(n-k+1,m)$
$=\frac{1}{q^{\ell}}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\left(\Psi(n-k+1,m)-\Psi(n-k,m)\right)$
$=\frac{1}{q^{\ell-1}(q-1)}\Psi(n-k,m)-\frac{1}{q^{\ell}(q-1)}\Psi(n-k+1,m)$
$=\frac{q^{n-\ell-k+1}}{q-1}\left(\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right)$
by (3).
We have
$\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)$ $=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+\frac{\rho^{(2)}(\xi)}{2!}\frac{1}{m^{2}}$
for some $\xi\in((n-k)/m,(n-k+1)/m)$, thus by Lemmas 12 and 13 we have
$\rho\left(\frac{n-k}{m}\right)-\rho\left(\frac{n-k+1}{m}\right)$ $=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+O\left(\frac{1}{m^{2}}\rho^{(2)}\left(\frac{n-k+1}{m}\right)\right)$
$=-\frac{1}{m}\rho^{\prime}\left(\frac{n-k+1}{m}\right)\left(1+O\left(\frac{\log(u+1)}{m}\right)\right).$
Then the result follows from Lemma 19. ∎
5.2. Minor arc estimates
Lemma 21.
For $\alpha,\beta\in\mathbf{T}$, we have
#$\#\left\{f\in{\mathcal{M}}(n):\|\alpha f-\beta\|<q^{-k}\right\}\leq\max\left\{1,q^{n-k},q^{n}\|\alpha\|,q^{-k}/\|\alpha\|\right\}.$
Proof.
Write
$\alpha=\sum_{i\leq 0}\alpha_{i}t^{i}\quad\text{and}\quad\beta=\sum_{i\leq 0}\beta_{i}t^{i}$
and let $\|\alpha\|=q^{-r}$.
If $\|\alpha f-\beta\|\leq q^{-k}$, then
(25) $\sum_{i=-j+r}^{n}f_{i}\alpha_{-j-i}=\beta_{-j},\quad\text{for }j=\max\{1,r-n+1\},\dots,\min\{k,r\}$
As $\alpha_{-r}\neq 0$, the coefficient matrix has full rank, and thus (25) has
$q^{n-(\min\{k,r\}-\max\{0,r-n\})}$
solutions. ∎
Lemma 22.
Let $\xi=a/g+\gamma$ with $\gcd(a,g)=1$, $\|\gamma\|<1/|g|^{2}$. Then we have
$S(\xi;n,m)\ll\left\{\begin{array}[]{c}m^{2}n^{1/2}\left(\frac{q^{n}}{|g|^{1/2}}+q^{3n/4+m/4}+|g|^{1/2}q^{(n+1)/2}\right),\\ m^{2}n^{1/2}\left(|g|^{1/2}q^{3n/4+m/4+1/2}+\|g\gamma\|^{1/2}q^{n}+\frac{q^{(n+1)/2}}{\|\gamma\|^{1/2}}\right).\end{array}\right.$
Proof.
[…]
$\Omega^{+}(u)\preccurlyeq\Omega^{-}(v),\quad\frac{n-m}{2}\leq\deg(u)<\frac{n+m}{2},\quad\text{and }\deg u+\deg\Omega^{-}(v)\geq\frac{n+m}{2}.$
For an irreducible polynomial $\omega\in{\mathcal{P}}$ write
${\mathcal{U}}(j,\omega)=\left\{u\in{\mathcal{S}}(j,m):\,\Omega^{+}(u)=\omega\right\}$
and
${\mathcal{V}}(j,\omega)=\left\{v\in{\mathcal{S}}(n-j,m):\,\Omega^{-}(v)\geq\omega,\deg\Omega^{-}(v)\geq\frac{n+m}{2}-j\right\}.$
Then, we have
$S\left(\xi;n,m\right)$ $=\sum_{\frac{n-m}{2}\leq j<\frac{n+m}{2}}\sum_{r=1}^{m}\sum_{\omega\in{\mathcal{P}}(r)}\sum_{u\in{\mathcal{U}}(j,\omega)}\sum_{v\in{\mathcal{V}}(j,\omega)}\mathbf{e}\left(uv\xi\right).$
Write
$S_{j,r}=\sum_{\omega\in{\mathcal{P}}(r)}\sum_{u\in{\mathcal{U}}(j,\omega)}\sum_{v\in{\mathcal{V}}(j,\omega)}\mathbf{e}\left(uv\xi\right).$
By the Cauchy inequality, we have
$|S_{j,r}|^{2}$ $\ll\left(\sum_{\omega\in{\mathcal{P}}(r)}\sum_{u\in{\mathcal{U}}(j,\omega)}1\right)\sum_{\omega\in{\mathcal{P}}(r)}\sum_{u\in{\mathcal{U}}(j,\omega)}\left|\sum_{v\in{\mathcal{V}}(j,\omega)}\mathbf{e}\left(uv\xi\right)\right|^{2}$
$\ll q^{j}\max_{\omega\in{\mathcal{P}}(r)}\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\left|\sum_{v\in{\mathcal{V}}(j,\omega)}\mathbf{e}\left(zv\xi\right)\right|^{2}$
$\ll q^{j}\max_{\omega\in{\mathcal{P}}(r)}\sum_{v_{1},v_{2}\in{\mathcal{V}}(j,\omega)}\left|\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\mathbf{e}(z(v_{1}-v_{2})\xi)\right|$
$\ll q^{j}\sum_{v_{1},v_{2}\in{\mathcal{P}}(n-j)}\left|\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\mathbf{e}(z(v_{1}-v_{2})\xi)\right|$
$\ll q^{j}\sum_{b\in\mathbb{F}_{q}^{*}}\sum_{v\in{\mathcal{P}}(n-j)}\sum_{i=0}^{n-j-1}\sum_{h\in{\mathcal{M}}(i)}\left|\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\mathbf{e}\left(bzh\xi\right)\right|+q^{n+j}.$
To prove the first bound, we apply Lemma 11 to get
$|S_{j,r}|^{2}$ $\ll q^{j}\sum_{b\in\mathbb{F}_{q}^{*}}\sum_{v\in{\mathcal{P}}(n-j)}\sum_{i=0}^{n-j-1}\left(\frac{q^{i+j}}{|g|}+q^{i}+|g|\right)+q^{n+j}$
$\ll n\left(\frac{q^{2n}}{|g|}+q^{2n-j}+|g|q^{n+1}+q^{n+j}\right).$
Whence
$S(\xi;n,m)\ll m^{2}n^{1/2}\left(\frac{q^{n}}{|g|^{1/2}}+q^{3n/4+m/4}+|g|^{1/2}q^{(n+1)/2}\right).$
To prove the second bound, we have by Lemma 7,
(26) $\sum_{h\in{\mathcal{M}}(i)}\left|\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\mathbf{e}\left(bzh\xi\right)\right|=q^{j}\cdot\#\left\{h\in{\mathcal{M}}(i):\|h\xi\|<q^{-j}\right\}.$
Write $h$ as $h=eg+f$, $|f|<|g|$. Then
$\left\|h\xi\right\|=\left\|eg\gamma+\frac{af}{g}+f\gamma\right\|.$
Putting $\beta=af/g+f\gamma$, then by Lemma 21,
#$\#\left\{|e|=\frac{q^{i}}{|g|}:\|eg\gamma+\beta\|<q^{-j}\right\}\ll\max\left\{1,\frac{q^{i-j}}{|g|},\|g\gamma\|\frac{q^{i}}{|g|},\frac{1}{\|g\gamma\|q^{j}}\right\}$
thus by (26)
$\sum_{h\in{\mathcal{M}}(i)}\left|\sum_{\begin{subarray}{c}z\in{\mathcal{M}}(j)\end{subarray}}\mathbf{e}\left(bzh\xi\right)\right|\ll|g|q^{j}+q^{i}+\|g\gamma\|q^{i+j}+\frac{1}{\|\gamma\|}.$
Whence
$|S_{j,r}|^{2}$ $\ll q^{j}\sum_{b\in\mathbb{F}_{q}^{*}}\sum_{v\in{\mathcal{P}}(n-j)}\sum_{i=0}^{n-j-1}\left(|g|q^{j}+q^{i}+\|g\gamma\|q^{i+j}+\frac{1}{\|\gamma\|}\right)+q^{n+j}$
$\ll n\left(|g|q^{n+j+1}+q^{2n-j}+\|g\gamma\|q^{2n}+\frac{q^{n+1}}{\|\gamma\|}\right).$
It yields
$S(\xi;n,m)\ll m^{2}n^{1/2}\left(|g|^{1/2}q^{3n/4+m/4+1/2}+\|g\gamma\|^{1/2}q^{n}+\frac{q^{(n+1)/2}}{\|\gamma\|^{1/2}}\right)$
6. Analysis of $S_{\mathcal{J}}$
We have the following results, for their proof see the proof of [12, Section 2.1].
Lemma 23.
We have
$\int_{\mathbf{T}}|S_{{\mathcal{J}}}(\xi)|\mathrm{d}\xi=1.$
Lemma 24.
We have
$S_{{\mathcal{J}}}(\xi)=\left\{\begin{array}[]{ll}q^{n-\#{\mathcal{I}}}\mathbf{e}(\xi t^{n})\prod_{i\in{\mathcal{I}}}\mathbf{e}(a_{i}t^{i}\xi)&\text{if }\xi_{-i-1}=0\text{ for all }i\not\in{\mathcal{I}},\\ 0&\text{otherwise.}\end{array}\right.$
The following result is [12, Lemma 2.7]
Lemma 25.
Let $a,g\in\mathbb{F}_{q}[t]$ be two polynomials with $\gcd(a,g)=1$ and write $g=g_{0}t^{k}$, $\gcd(g_{0},t)=1$. If $11<|g_{0}|\leq q^{\lceil n/(\#{\mathcal{I}}+1)\rceil}-$ 1, then
${\mathcal{S}}_{{\mathcal{J}}}\left(\frac{a}{g}\right)=0.$
The following result is the analogue of Bourgain’s result [2, Lemma 3] proved by Ha [12, Lemma 2.6]
Lemma 26.
Let $k\leq n/2$. Then we have
$\sum_{\begin{subarray}{c}|g|<q^{k}\\ \gcd(aT,g)=1\end{subarray}}\left|S_{\mathcal{J}}\left(\frac{a}{g}\right)\right|\leq\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2kC_{1}\#{\mathcal{I}}/n},$
where $C_{1}=\kappa(n,2k,\#{\mathcal{I}}/n)\leq 2$ is given by
$\kappa(x,y,\delta)=\left\{\begin{array}[]{cl}\frac{2}{\delta+1}&\text{if }1<x/y<2\\ \frac{2v}{(v+1)\delta+v-1}&\text{if }v-1<x/y<v\\ 1&\text{if }y\mid x.\end{array}\right.$
7. Proof of Theorem 3
First we investigate the integral (9) over the major arcs defined by (10). We also choose
(27) $\kappa=\left\lceil\frac{n-m-2}{4}\right\rceil\quad\text{and}\quad\ell=\left\lceil\frac{n-m}{4}\right\rceil.$
As the polynomials $f$ in the summations $S(\xi;n,m)$ and $S_{\mathcal{J}}(\xi)$ are monic and of degree $n$, we have for $\xi\in{\mathfrak{M}}$ having the form $\xi=a/g+\gamma$, $|\gamma|<q^{-n+\ell}$ that $\mathbf{e}(f\gamma)=\mathbf{e}(b\gamma)$ if $f\equiv b\mod{\mathcal{R}}_{\ell}$, $b\in{\mathcal{M}}(n)$. Then
$\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi$
$=\sum_{\begin{subarray}{c}a,g\\ |g|\leq q^{\kappa}\end{subarray}}\int_{|\gamma|<q^{-n+\ell}}S\left(\frac{a}{g}+\gamma;n,m\right)\overline{S_{\mathcal{J}}\left(\frac{a}{g}+\gamma\right)}\mathrm{d}\gamma$
$=\sum_{\begin{subarray}{c}a,g\\ |g|\leq q^{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g}\right)\right)\int_{|\gamma|<q^{-n+\ell}}\mathbf{e}\left(b\gamma\right)\overline{\mathbf{e}\left(b\gamma\right)}\mathrm{d}\gamma$
$=\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}a,g\\ |g|\leq q^{\kappa}\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g}\right)\right).$
We split the sum above into two parts
(28) $\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi=M_{1}+M_{2}$
with
$M_{1}=\frac{1}{q^{n-\ell}}\sum_{0\leq k\leq\kappa}\sum_{\begin{subarray}{c}|a|<t^{k}\\ t\nmid a\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{t^{k}}\right)\right)$
and
$M_{2}=\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}a,g\\ g\neq t^{k}\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g}\right)\right).$
Concerning $M_{1}$, for $k=0$ we have by Lemma 18,
(29) $\begin{split}&\frac{1}{q^{n-\ell}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\right)\\ =&\frac{1}{q^{n-\ell}}\left(\frac{1}{q^{\ell}}\Psi(n,m)+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n}e^{\varepsilon\ell}\right)\right)\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}1\\ =&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right).\end{split}$
Now consider the terms with $g=t^{k}$, $k\geq 1$. Let $B={\mathcal{I}}\cap\{n-\ell,\dots,n-1\}$. Clearly, one can represent the residue classes of ${\mathcal{R}}_{\ell}$ by the elements of ${\mathcal{M}}(\ell)$. The contribution of those $b$ that $b_{i-(n-\ell)}\neq\alpha_{i}$ for $i\in B$ is zero. Write
=${\mathcal{B}}=\left\{b\in{\mathcal{M}}(\ell):b_{j-(n-\ell)}=\alpha_{j}\text{ for }j\in B\right\}\quad\text{and}\quad I_{\ell}=I\cup\{n-\ell,\dots,n-1\}$
and for $b\in{\mathcal{B}}$ write
${\mathcal{J}}_{b}=\left\{f\in{\mathcal{M}}(n):f\in{\mathcal{J}}\text{ and }f\equiv b\mod{\mathcal{R}}_{\ell}\right\}$
and
$\beta_{i}=\left\{\begin{array}[]{cl}\alpha_{i}&\text{if $i\in{\mathcal{I}}$},\\ b_{i-(n-\ell)}&\text{if $i\in\{n-\ell,\dots,n-1\}\setminus I$ }.\end{array}\right.$
Then
#$\#{\mathcal{J}}_{b}=q^{n-\#{\mathcal{I}}_{\ell}}=q^{n-\#{\mathcal{I}}-\ell+\#B}$
and
$S_{{\mathcal{J}}_{b}}(\xi)=\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(f\xi\right).$
For $1\leq k\leq\kappa$ we recall that by Lemmas 12, 13 and 20, we have
$\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)=-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+E(n,m,k)$
with error term
$E(n,m,k)\ll q^{n-\ell-k}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m^{2}}+q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{k}{2}}e^{\varepsilon k}.$
By Lemma 24, $S_{{\mathcal{J}}_{b}}(a/t^{k})=0$ if $a_{i}\neq 0$ for some $k-i-1\not\in{\mathcal{I}}_{\ell}$, $0\leq i<k$. As $\gcd(a,t)=1$, we have $a_{0}\neq 0$ and thus it is enough to consider $k$ with $k-1\in{\mathcal{I}}_{\ell}$. For such a $k$, we have by Lemma 24 that
(30) $\begin{split}&\frac{1}{q^{n-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{t^{k}}\right)S_{{\mathcal{J}}_{b}}(a/t^{k})\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\sum_{\begin{subarray}{c}f\in S(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}\frac{a_{i}t^{i}f}{t^{k}}\right)\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}_{\ell}-\ell}}\left(-\frac{1}{m}\frac{q^{n-\ell-k+1}}{q-1}\rho^{\prime}\left(\frac{n-k+1}{m}\right)+E(n,m,k)\right)\cdot S_{I}(n,k,\ell)\end{split}$
with
$S_{I}(n,k,\ell)=\sum_{b\in{\mathcal{B}}}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}_{\ell}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}_{\ell}}\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right).$
Considering $S_{I}(n,k,\ell)$, as $k-i-1\leq k\leq\kappa<n-\ell$, for the summation over $a_{i}$ it is enough to consider $i$ with $k-i-1\in{\mathcal{I}}$. Moreover, for $j\geq n-\ell$ and $0\leq i<k$, we have $j+i-k\geq n-\ell-k>0$, and thus
$\mathbf{e}\left(\beta_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}_{\ell}\end{subarray}}a_{i}t^{i-k}\right)\right)=1,$
i.e., the terms in $S_{I}(n,k,\ell)$ do not depend on the choice of $b\in{\mathcal{B}}$. Therefore
$S_{I}(n,k,\ell)$ $=q^{\ell-\#B}\sum_{\begin{subarray}{c}a_{i}\in\mathbb{F}_{q}\\ k-i-1\in{\mathcal{I}}\\ 0\leq i<k\\ a_{0}\neq 0\end{subarray}}\prod_{j\in{\mathcal{I}}}\mathbf{e}\left(\alpha_{j}t^{j}\left(\sum_{\begin{subarray}{c}0\leq i<k\\ k-i-1\in{\mathcal{I}}\end{subarray}}a_{i}t^{i-k}\right)\right)$
$=q^{\ell-\#B}\prod_{\begin{subarray}{c}j\in{\mathcal{I}}\\ 0\leq j<k-1\end{subarray}}\left(\sum_{a_{k-j-1}\in\mathbb{F}_{q}}\mathbf{e}\left(\alpha_{j}a_{k-j-1}t^{-1}\right)\right)\left(\sum_{a_{0}\in\mathbb{F}_{q}^{*}}\mathbf{e}\left(\alpha_{k-1}a_{0}t^{-1}\right)\right)$
$=\sigma_{k}q^{\ell-\#B+\#{\mathcal{I}}\cap\{0,\dots,k-2\}}\mathbf{1}(\alpha_{j}=0,j\in{\mathcal{I}}\cap\{0,\dots,k-2\}),$
where
$\sigma_{k}=\left\{\begin{array}[]{cl}-1&\alpha_{k-1}\neq 0,\\ q-1&\text{otherwise}.\end{array}\right.$
Write $I=\{i_{0}<i_{1}<\dots<i_{\nu}<\dots\}$ with $a_{i_{0}}=a_{i_{1}}=\dots=a_{i_{\nu-1}}=0\neq a_{i_{\nu}}$. Then taking summation over $0\leq k\leq\kappa$ we get by (29) and (30), that
(31) $\begin{split}M_{1}=&\frac{\Psi(n,m)}{q^{\#{\mathcal{I}}}}+O\left(q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\ell}e^{\varepsilon\ell}\right)\\ &+(q-1)\sum_{\begin{subarray}{c}0\leq j<\nu\\ i_{j}<\kappa\end{subarray}}\frac{1}{q^{\#{\mathcal{I}}-j-\ell}}\left(-\frac{1}{m}\frac{q^{n-i_{j}-\ell}}{q-1}\rho^{\prime}\left(\frac{n-i_{j}}{m}\right)+E(n,m,i_{j}+1)\right)\\ &-\mathbf{1}(\nu<\kappa)\frac{1}{q^{\#{\mathcal{I}}-\nu-\ell}}\left(-\frac{1}{m}\frac{q^{n-i_{\nu}-\ell}}{q-1}\rho^{\prime}\left(\frac{n-i_{\nu}-\ell}{m}\right)+E(n,m,i_{\nu}+1)\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)\right)\\ &+O\left(\nu q^{n-\#{\mathcal{I}}-i_{0}-1}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m^{2}}+q^{\left(\frac{1}{2}+\varepsilon\right)n-\#{\mathcal{I}}+\frac{\kappa}{2}+\ell}e^{\varepsilon(n-m)/2}\right)\\ =&\frac{1}{q^{\#{\mathcal{I}}}}\Psi(n,m)+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)\right)\\ &+O\left(q^{n-\#{\mathcal{I}}-i_{0}-1}\rho\left(u\right)\frac{u^{1/2}(\log(u+1))^{5/2}}{m}+q^{\left(\frac{7}{8}+\varepsilon\right)n-\frac{3m}{8}+\frac{5}{4}-\#{\mathcal{I}}}e^{\varepsilon(n-m)/2}\right)\end{split}$
by (27).
For $M_{2}$, write $K=\lceil n/(\#{\mathcal{I}}+1)\rceil$. If $g$ has the form $g=g_{0}t^{k}$ with $t\nmid g_{0}$ and $1<|g_{0}|<q^{K}$, then $S_{{\mathcal{J}}}(a/g)$ vanishes by Lemma 25. Then, we can write
$M_{2}=\frac{1}{q^{n-\ell}}\sum_{r=K}^{\kappa}\sum_{k=0}^{\kappa-r}M_{2}(r,k)$
with
$M_{2}(r,k)=\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{\begin{subarray}{c}f\in{\mathcal{S}}(n,m)\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(\frac{af}{g_{0}t^{k}}\right)\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)\right).$
By Lemma 19,
$\begin{split}M_{2}(r,k)&=\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\left(\sum_{d\mid g_{0}t^{k}}\frac{\mu(g_{0}t^{k}/d)}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)\end{split}$
with error term
(32) $R(r,k)\ll q^{\left(\frac{1}{2}+\varepsilon\right)n+\frac{r+k}{2}}e^{\varepsilon(\ell+\kappa)}.$
By (12) and Lemmas 12 and 15, we have
$\sum_{d\mid g_{0}t^{k}}\frac{\mu(g_{0}t^{k}/d)}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi_{g_{0}t^{k}/d}(n-\deg d,m)$
$\leq$ $\sum_{d\mid g_{0}t^{k}}\frac{1}{q^{\ell}\Phi(g_{0}t^{k}/d)}\Psi(n-\deg d,m)$
$\ll$ $\sum_{d\mid g_{0}t^{k}}\frac{q^{n-\deg d}}{q^{\ell}\Phi(g_{0}t^{k}/d)}\rho\left(\frac{n-\deg d}{m}\right)$
$\ll$ $q^{n-\ell}\rho\left(u\right)\sum_{d\mid g_{0}t^{k}}\frac{(u\log u)^{\deg d/m}}{|d|\Phi(g_{0}t^{k}/d)}$
$\ll$ $q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}(r+k)\left(1+\log_{q}(r+k)\right).$
Whence
(33) $\begin{split}M_{2}(r,k)&=\left(q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}\kappa\left(1+\log_{q}\kappa\right)+R(r,k)\right)\\ &\qquad\cdot\sum_{\begin{subarray}{c}\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right).\end{split}$
In order to handle the second term, write
$\frac{af}{g_{0}t^{k}}=\frac{a_{1}f}{g_{0}}+\frac{a_{2}f}{t^{k}}\quad\text{with}\quad\gcd(a_{1},g_{0})=\gcd(a_{2},t)=1.$
Then we have by Lemmas 9 and 26 that
$\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\sum_{\begin{subarray}{c}b\bmod{\mathcal{R}}_{\ell}\\ b\in{\mathcal{M}}(n)\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv b\bmod{\mathcal{R}}_{\ell}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ t\nmid g_{0}\end{subarray}}\sum_{\begin{subarray}{c}|a|<|g_{0}t^{k}|\\ \gcd(a,g_{0}t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\end{subarray}}\mathbf{e}\left(-\frac{af}{g_{0}t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{t},t)=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}-\frac{a_{2}f}{t^{k}}\right)$
$=$ $\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{2}|<|t^{k}|\\ \gcd(a_{t},t)=1\end{subarray}}\mathbf{e}\left(-\frac{a_{2}c}{t^{k}}\right)\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)$
$=$ $\sum_{\begin{subarray}{c}|c|<|t^{k}|\\ \gcd(c,t)=1\end{subarray}}\mu(t^{k})\sum_{\begin{subarray}{c}g_{0}\in{\mathcal{M}}(r)\\ \gcd(g_{0},t)=1\end{subarray}}\sum_{\begin{subarray}{c}|a_{1}|<|g_{0}|\\ \gcd(a_{1},g_{0})=1\end{subarray}}\sum_{\begin{subarray}{c}f\in{\mathcal{J}}\\ f\equiv c\bmod t^{k}\end{subarray}}\mathbf{e}\left(-\frac{a_{1}f}{g_{0}}\right)$
$\ll$ $|\mu(t^{k})|\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+k)/n}.$
where $C_{1}$ as in Lemma 26. Thus by (33)
(34) $\begin{split}M_{2}(r,k)&\ll\left(q^{n-\ell-r-k}\rho\left(u\right)\left(u\log u\right)^{1/8}(r+k)\left(1+\log_{q}(r+k)\right)+R(r,k)\right)\\ &\quad\cdot|\mu(t^{k})|\frac{q^{n}}{q^{\#{\mathcal{I}}}}q^{2rC_{1}(\#{\mathcal{I}}+k)/n}.\end{split}$
Write
$C_{2}=1-4C_{1}\frac{\#{\mathcal{I}}+1}{n}.$
As #$\#{\mathcal{I}}<n/10$, we have then $C_{2}\geq 1/5$ and $22C_{1}(\#I+1)/n\leq 1/$ 2. As $M_{2}(r,k)=0$ for $k\geq 2$, we have by (32) and (34) that
(35) $\begin{split}M_{2}&\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\sum_{r\geq K}\frac{r\left(1+\log_{q}r\right)}{q^{C_{2}r}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right)\\ &\ll\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\rho\left(u\right)\left(u\log u\right)^{1/8}\frac{K\left(1+\log_{q}K\right)}{q^{C_{2}K}}+O\left(q^{-\frac{(1-4\varepsilon)n+m}{4}+2}e^{\varepsilon(n-m)/4}\right)\right).\end{split}$
Then, combining (28), (31) and (35), we get that the contribution of the major arcs is
(36) $\begin{split}\int_{{\mathfrak{M}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&=\frac{\Psi(n,m)}{q^{\#I}}\left(1+O\left(\left(u\log u\right)^{1/8}\frac{\delta^{-1}\log\delta}{q^{C/\delta}}+\frac{1}{q^{i_{0}+1}}\frac{u^{1/2}(\log u)^{5/2}}{m}\right)\right)\\ &\quad+\frac{q^{n}}{q^{\#{\mathcal{I}}}}\left(\Lambda_{0}(n,m,I,u)+\Lambda_{1}(n,m,I,u)+O\left(q^{-\frac{(1-8\varepsilon)n+2m}{8}+2}e^{\varepsilon\frac{n-m}{2}}\right)\right).\end{split}$
Next, we consider the integral along the minor arcs ${\mathfrak{m}}$
$\int_{{\mathfrak{m}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi.$
By Lemma 14, write $\xi\in\mathbf{T}$ as
$\xi=\frac{a}{g}+\gamma,\quad\gcd(a,g)=1,|g|<q^{n/2},|\gamma|\leq\frac{1}{|g|q^{n/2}}.$
If $\xi\in{\mathfrak{m}}$, then either $q^{\kappa}\leq|g|<q^{n/2}$ or $|g|<q^{\kappa}$ and $q^{-n+\ell}\leq|\gamma|\leq 1/(|g|q^{n/2})$. In the formal case, we estimate $S(\xi;n,m)$ by the first bound of Lemma 22
$S(\xi,m,n)$ $\ll m^{2}n^{1/2}\left(\frac{q^{n}}{|g|^{1/2}}+q^{3n/4+m/4}+|g|^{1/2}q^{(n+1)/2}\right)$
$\ll m^{2}n^{1/2}\left(q^{n-\kappa/2}+q^{3n/4+m/4}\right).$
In the second case, we use the second bound of Lemma 22
$S(\xi,m,n)$ $\ll m^{2}n^{1/2}\left(|g|^{1/2}q^{3n/4+m/4+1/2}+\|g\gamma\|^{1/2}q^{n}+\frac{q^{(n+1)/2}}{\|\gamma\|^{1/2}}\right)$
$\ll m^{2}n^{1/2}\left(q^{3n/4+m/4+\kappa/2+1/2}+q^{n-(\ell-1)/2}\right).$
By (27), we get in both cases
$S(\xi,m,n)\ll m^{2}n^{1/2}q^{(7n+m+4)/8}.$
Thus by Lemma 23, we have
(37) $\begin{split}\int_{{\mathfrak{m}}}S(\xi;n,m)\overline{S_{\mathcal{J}}(\xi)}\mathrm{d}\xi&\ll m^{2}n^{1/2}q^{(7n+m+4)/8}\int_{{\mathfrak{m}}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{(7n+m+4)/8}\int_{\mathbf{T}}|S_{\mathcal{J}}(\xi)|\mathrm{d}\xi\\ &\ll m^{2}n^{1/2}q^{(7n+m+4)/8}.\end{split}$
By replacing $\varepsilon$ if needed in (36), Theorem 3 follows from (36)and (37).
8. Proof of Theorem 1
Theorem 1 follows directly form Theorem 3. Indeed, if $0\in{\mathcal{I}}$ and $\alpha_{0}\neq 0$, then $\Lambda_{0}(n,m,I,u)=0$ and
$\Lambda_{1}(n,m,I,u)=-\frac{1}{m}\frac{1}{q-1}\rho^{\prime}\left(u\right)\sim\frac{1}{m}\frac{1}{q}\rho\left(u\right)\log u$
by Lemmas 12 and 13, which is absorbed by the error term.
9. Proof of Corollaries 2 and 4
Corollary 2 follows directly from Theorem 1.
For Corollary 4, if $q\rightarrow\infty$, then
$|\Lambda_{0}(n,m,I,u)|,|\Lambda_{1}(n,m,I,u)|\leq c(n,m)\frac{1}{q}$
for some constant $c(n,m)$ depends only on $n$ and $m$ (and thus on $u$).
If $n\rightarrow\infty$ , then by Lemmas 12 and 13,
$|\rho^{\prime}((n-i_{j})/m)|\leq c\rho(u)u^{1/8}(\log u)^{9/8}$
for some absolute constant $c$, and as $\nu\leq\#{\mathcal{I}}=n\delta$, thus
$|\Lambda_{0}(n,m,I,u)|\leq c(q)\rho(u)\frac{u^{1/8}(\log u)^{9/8}}{m}n\delta=c(q)\rho(u)u^{9/8}(\log u)^{9/8}\delta=o(\rho(u)).$
Similarly,
$|\Lambda_{1}(n,m,I,u)|\leq c_{1}(q)\frac{1}{m}\rho^{\prime}\left(\frac{n-i_{\nu}}{m}\right)\leq c_{2}(q)\rho(u)\frac{u^{1/8}(\log u)^{9/8}}{m}=o(\rho(u)).$
Acknowledgment
The author wishes to thank Ofir Gorodetsky and Eugenijus Manstavic̆ius for useful discussions and valuable comments. He also thanks the anonymous referee to bring the attention to [1] which allowed to improve Lemma 16.
The author was was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960,